Currington Company wants to use absorption cost-plus pricing to set the selling price on a newly remodeled product. The company plans to invest $158,000 in operating assets to produce and sell 12,000 units. Its required return on investment (ROI) in its operating assets is 16%. The accounting department has provided cost estimates for the new product as follows:

Per Unit Total
Direct materials $ 4.40
Direct labor $ 3.40
Variable manufacturing overhead $ 1.40
Fixed manufacturing overhead $ 70,800
Variable selling and administrative expenses $ 1.40
Fixed selling and administrative expenses $ 12,280

Required:

1. What is the unit product cost for the remodeled product? (Round intermediate calculations and final answer to 2 decimal places.)

2. What is the markup percentage on the absorption cost for the remodeled product? (Round intermediate calculations to 2 decimal places.)

3. What selling price would the company establish for its remolded product using a markup percentage on absorption cost? (Round intermediate calculations and final answer to 2 decimal places.)

4. Suppose the company actually produced and sold only 10,000 units (instead of its planned sales volume of 12,000 units) at the selling price that you derived in requirement 3. What ROI did the company actually earn at this lower sales volume? (Round intermediate calculations to 2 decimal places. Round your percentage answer to 1 decimal place.)

5. Assume that the company wants to raise the price of its newly remodeled product with the intention of achieving the product’s desired ROI at the lower sales volume of 10,000 units. Using absorption cost-plus pricing, what would be the revised selling price at this lower sales volume? (Round your intermediate percentages to 1 decimal place. Round all other intermediate calculations to 2 decimal places.)

1. Unit product cost
2. Markup percentage on absorption cost %
3. Selling price per unit
4. Return on investment %
5. Revised selling price per unit
Answer #1

1.Unit product cost of remodeled product = Direct material + Direct labor + Variable manufacturing overhead + Fixed manufacturing overhead

= 4.40+3.40+1.40+70,800/12,000

= $15.1 per unit

2.Desired Return = 158000*16% = $25,280

Mark up % = (Desired Return+ Selling and Admin Expenses)/Product Cost

= (25,280 + 1.40*12000+12,280)/(15.1*12000)

= 54,360/181,200

= 30%

3.Selling price = Product cost + Markup

= 15.1 + 15.1*30%

= $19.63 per unit

4.ROI = Operating Income/Investment

=[(19.63 -4.40-3.40-1.40-1.40)*10,000 – 70,800-12,280]/158000

= 4.5696%

i.e. 4.57%

5.Revised Cost per Unit = 4.40+3.40+1.40+70,800/10,000

= $16.28 per unit

Mark up = (25,280 + 1.40*10000+12,280)/(16.28*10000)

= 31.67%

Revised Price = $21.435876

i.e. $21.44 per unit